A compound that contains only carbon, hydrogen, and oxygen is 45.3 % C and 9.43 % H by mass.
What is the empirical formula of the substance?

Enter the correct subscript for each atom using the smallest whole number.


Its not C1H3O1
Answer #1

mass C in 100 g = 45.3% of 100 g = 45.3 g
moles C = mass / molar mass = 45.3 g / 12.01 g/mol = 3.77 moles

mass H in 100 g = 9.43 g
moles H = 9.43 g / 1.0 g/mol = 9.43 mol

mass O in 100 g = 45.27 g
moles O = 45.27 g / 16.00 g/mol = 2.829 mol

ratio of moles C : H : O
= 3.77 : 9.43 : 2.829

divide each number by samllest number in ratio (2.829)

C : H : O
= 1.33 : 3.33 : 1.0

This is not a whole number ratio yet, multiply by the smallest number needed to get all whole number (in this case x3)

3 x (1.33 : 3.33 : 1.0)
= 4 : 10 : 3

C4H10O3
Answer #2

C4H10O3